# Pragmatic Inference for Moral Reasoning Acquisition: Generalization via Distributional Semantics

### v2

@@ 1 Introduction
With the advancement of Large Language Models (LLMs), how to enhance their capability of moral reasoning in order to regulate their performance in accordance with human moral values has attracted scholarly interest. However, it still remains as a challenging task due to the gap between distributional semantics used by LLMs and the pragmatics-based nature of moral reasoning  liu2025diagnosing; metaphorically speaking, the gap between what is said and what is morally implied. Distributional semantics follow Firth’s (firth1957synopsis) distributional hypothesis which argues that linguistic units have similar meaning if they occur in similar contexts. In contrast, moral reasoning goes beyond this structuralism view of language and requires the capability of deriving conclusions from implicature of language use, moral norms and behavioral consequences, and observations of discursive rules of conduct spencer-oatey_interdisciplinary_2019. These unsaid variables need to be linguistically manifested so that the gap between distributional semantics and pragmatics in moral reasoning can be closed, which increases the methodological generalizability that previous studies have thus far struggled to achieve rao2023ethical; preniqi2024moralbert; jiang2025investigating.
@@
Table 1: Moral Reasoning Example From ziems2022moral. The Moral Judgment for the given reply is Agree, and the Rule of Thumb (RoT) represents a judgment regarding an action mentioned in the prompt-reply, grounded in the moral foundation of fairness.
Table 1 demonstrates an example of the several elements that are involved in moral reasoning: rule of thumb (RoT), moral foundation, and moral judgment ziems2022moral. Thus far, learning paradigms, which aim to equip LLMs with moral reasoning, generally train models to directly map a situation (Prompt and Reply) to moral judgment, with or without a reference to moral foundations and/or RoT  hendrycks2020aligning; forbes2020social; jiang2021can; ziems2022moral; jiang2025investigating. Direct mapping often requires a huge amount of data to enable LLMs to capture the statistical correlations between situations and moral reasoning objectives.
On the other hand, we have witnessed the effectiveness of pragmatic inference used in other morality-relevant tasks, such as the identification of social biases and toxicity  sap2020social; zhou-etal-2023-cobra; chen2025pragmatic. For example,  sap2020social developed a structured inference path that enables the identification of stereotypical implications encoded in explicit toxic language. chen2025pragmatic developed an inference chain that provides LLMs with the freedom to interpret online neologisms and rephrase the literal meanings of an input. Their method improved the performance of LLMs in recognizing highly implicit toxic language by 12% to 20%. These approaches are consonant with our goal of bridging the gap between distributional semantics and pragmatics, for instance, through interpreting literal meanings chen2025pragmatic.
However, moral reasoning requires different inferential methods from the aforementioned toxicity or bias identification. Unlike toxic language (e.g., “lowering employment standards for females” implying “females are less capable”), the moral foundations***Please refer to Section 2 for more discussion on Moral Foundations Theory for moral reasoning. that underlie a moral situation are often not encoded by certain linguistic cues at all  clark1989contributing; graesser1994constructing; sap-etal-2022-neural. Inferring them necessitates context-specific interpretations of abstract moral rules (e.g., the rule “do not hurt others” may manifest as “do not push another child to the floor” on a children’s playground and “do not emotionally abuse your partner” for adult relationships). Therefore, abstract moral foundations, rules of conduct, and context all need to be sufficiently textualized so as to be incorporated in machines’ reasoning process.
@@ 2 Related Works
Distributional Semantics is one of the theoretical foundations for embeddings. It interprets the meaning of a linguistic unit by additional signs (e.g., words, emoji) surrounding it harris1954distributional; firth1957synopsis; mcdonald2001testing; lenci2023distributional; boleda2020distributional. However, due to its nature as a structuralism approach to understanding language, distributional semantics suffers limitations. It treats the “context” purely as the texts provided, neglecting the many “unsaid” and latent parts of real interactions, such as the social constraints that a speaker encounters, the effects that a speech creates on the listeners, and the delivery of intention lenci2008distributional; purver2015distributional. These are studied in pragmatics, where LLMs have shown several deficiencies, e.g., understanding implicature cong_manner_2024, comprehending metaphors and humours ruis_goldilocks_2023; barattieri_di_sanpietro_pragmatic_2023, and moral alignment khamassi2024strong; liu2024towards.
@@
Pragmatic Inference and Moral Reasoning. In this work, we leverage pragmatic inference to facilitate LLMs to achieve stronger performance in moral reasoning. Pragmatic inference refers to the process of deriving conclusions of meanings based on context, intention, and language use elder_pragmatic_2024. The inferential process is often not verbally realized and, instead, involves evaluations made at the metapragmatic level kadar_understanding_2013. That is, humans construct meta links that connect an instance of language use to certain social qualities (e.g., linking the phrase “thank you” to the concept of politeness). They learn these metapragmatic links over their physical and social experiences, which are entirely absent in LLMs. These limitations suggest that: (i) the inferential process, that is not verbally realized and often does not need to be for humans, needs to be textualized so that LLMs can learn; and (ii) the context, where humans gain their experiences, needs to be sufficiently textualized, including both societal (e.g., social norms) and situational contexts (e.g., situated interpretations of language use). Without effective methods to textualize them, it is not a surprise that moral reasoning has long been a challenge in the broader field of artificial intelligence allen2005artificial; van2014hybrid; wallach2020machine.
@@
Moral Foundations Theory (MFT). haidt2004intuitive and graham2013moral describe a framework for understanding human morals and moral reasoning. MFT hypothesizes that moral judgments are guided by domain-specific foundations, i.e., care, fairness, loyalty, liberty, authority, and sanctity. These foundations capture the diversity of moral concerns and have been widely adopted in computational social science johnson2018classification; wendell2021classifying; reinig2024survey. They provide a principled way to annotate, analyze, and model moral dimensions in language, making them particularly relevant for studying how LLMs represent and reason about morality forbes2020social; ziems2022moral.
@@ 3 Preliminary
Benchmark. In this paper, we adopt the Moral Integration Corpus (MIC) benchmark and its annotations from ziems2022moral. We chose the MIC benchmark because it closely reflects how humans make moral judgments in daily communication and provides richer contextual information. One advantage of the MIC dataset is that it provides levels of agreement between the annotators, which allows us to focus on the less controversial part of their data. Table 1 presents an example from MIC. The example consists of a Prompt (“Do people actually shit themselves when they get very scared??”) and a Reply (“I don’t think I’ve ever heard of that happening, but I guess it’s possible..”), with three evaluations annotated: agreement with the reply as moral judgment, RoT, and MF. The MIC benchmark utilizes the six MFs defined by graham2013moral: care, fairness, liberty, loyalty, authority, and sanctity. We take two of the evaluations as our objectives: moral judgment (agree, neutral, or disagree with the reply) and MFs. RoT, instead, is used as part of the learning paradigm. RoT is not selected as a learning objective because: (i) it does not have a fixed answer and, more crucially, multiple RoTs may exist in one situationWe refer readers to Section 3.1 of ziems2022moral for further details., and (ii) it has a rather dynamic relationship with MFs.
Experimental Settings. Following the steps of our pragmatic inference frameworks (detailed in Section 4), we design several questions and prompt the deepseek model to answer them. The step-wise answers are then used as the training dataThe data used in our experiments is available in the anonymized GitHub repository: https://anonymous.4open.science/r/moralpragmatics-F891. for fine-tuning the LLMs towards moral reasoning objectives. We chose Llama3-3.2-1B, Llama3-3.2-3B grattafiori2024llama3herdmodels and Mistral-7B (Mistral-7B-v0.3 jiang2023mistral7b) as our backbone models, and leverage the supervised fine-tuning objective to train those models by following previous studies ziems2022moral; liu2025diagnosing. We intentionally select smaller models, as they are generally considered less capable than their larger counterparts. We aim to demonstrate the effectiveness of our pragmatic inference frameworks through the strong performance achieved by these smaller models.
@@
For baseline approaches, we follow prior work forbes2020social; ziems2022moral and consider two settings: (1) The base setting. For the moral judgment task, we compute $y_{J}=f_{\theta}(x_{S})$; and for MFC, we compute $y_{M}=f_{\theta}(x_{R})$. (2) The base+ setting. For the moral judgment task, we compute $y_{J}=f_{\theta}(x_{S},y_{M})$; and for MFC task, we compute $y_{M}=f_{\theta}(x_{R},d_{m})$. We additionally include the definitions of the six MFs ($d_{m}$) in the prompts. When the objective is moral judgment, we also explicitly provide the ground-truth MFs in the prompts. In other words, we intentionally provide the models with all available information in order to build strong baselines, which are then used to compare with the results of our framework. The setting of our proposed method is in line with the base+ setting but we incorporate an additional inference process in the input, e.g., $y_{J}=f_{\theta}(x_{S},y_{M},\{\text{inference}\})$.
@@ 4.1 Motivation
Proposing a pragmatic inference method for moral reasoning is challenging because there is no linguistic theory that clearly defines what constitutes an appropriate pragmatic inference for LLMs. Although several theories describe how humans make moral judgments, LLMs lack human-like linguistic competence bender-koller-2020-climbing; mahowald2024dissociating, making these theories unsuitable for direct application. On the other hand, LLMs are well known to capture distributional knowledge (i.e., distributional semantics) through statistical associations, and prior works have shown that they cannot leverage this distributional knowledge for satisfactory moral reasoning performance liu-etal-2025-diagnosing; lenci2023distributional. However, the distributional knowledge is much beyond distributional semantics, this is because pre-training corpora also contains static and context-limited texts graesser1994constructing; sap-etal-2022-neural, forcing LLMs not only capture distributional semantics but also approximate pragmatic cues through shallow statistical correlations.
For instance, the statement “calling your uncle by his given name is rude” is acceptable in many East Asian cultures but may not be perceived the same way elsewhere. Because pre-training texts provide no explicit contextual cues about the cultural setting in which such statements are embedded, LLMs may learn this norm only by associating negative descriptors “is rude” with the action “calling your uncle by his name” even though they lack any actual cultural understanding pawar2025survey. This implies that LLMs can leverage these superficial associations to predict certain pragmatic cues. However, the extent to which they can do so remains unknown, and we cannot rely on a purely semantics-driven approach to simplify the pragmatic inference. We hypothesize that MFT-driven pragmatic inference can help LLMs address the challenges outlined above, based on the following two perspectives.
@@
From the linguistic standpoint, our design of pragmatic inference methods is guided by three principles: (1) MFs are common ground which underlie all interactions. This has been supported by the theory and experiments of MFs graham2013moral; abdulhai2024moral. (2) MFs are abstract values extracted from discursive situations and hence need to be reified as context-specific interpretations graham_mapping_2011. It is these context-specific interpretations of MFs which provide effective semantic cues for the models to reason morally. (3) Moral judgment is the judgment of good or bad by its nature, more specifically, the good/bad evaluation placed on the outcome of an action nozick1968moral; schaich2006consequences; emelin2021moral. Therefore, it is important to build the connections between contexts, discursive interpretations of MFs, and the outcome (i.e., behavioral consequence).
From a statistical standpoint, liu2025diagnosing demonstrate that the pragmatic gap between moral situations and moral reasoning objectives introduces greater degradation in language modeling performance than purely semantics-driven tasks. This is because LLMs tend to develop spurious associations, between moral situations and reasoning objectives, that undermine distributional semantics. To address this, MFT provides a principled signal: all moral situations can be interpreted through its lens. The key, then, is to leverage our pragmatic inference methods to supply linguistic cues that appropriately map moral situations to the LLMs’ hidden state space that is characterized by MFs, thereby alleviating the degradation of distributional semantics.
@@ 4.2 Moral Foundations Classification (MFC)
Figure 1 presents the prompting questions of the inferential steps used in the MFC task. The input here is a simple RoT and the definitions of six MFs. The inference steps are built as such: (1) Step 1 identifies the action encoded by the RoT forbes2020social. This step prepares the target (action) for moral evaluation. (2) Step 2 elicits the outcome of the action, namely, the behavioral consequence. One element of pragmatic inference is social norms, the variable that is probably most important in this task, given that RoTs were provided here without context. Norms for moral evaluations feature the physical and emotional harm to others that an action may incur huebner_moral-conventional_2010; kadar_impoliteness_2015; Turiel1983-TURTDO. Therefore, Step 2 asks for the behavioral consequence as the linguistic manifestation of the social norms. This step setting also aligns with prior NLP studies showing that outcomes help establish the link between an action and its moral judgment lindner2017hera; lindner2019extracting. (3) Step 3 connects the action and its consequence to relevant MFs by textualizing their metapragmatic links.
Table 2 reports performance on development set instances with 1–3 moral foundation labels. Moral situations with more than three labels are relatively rare. We also report the average performance for each experimental setting. Across all training dataset sizes, model variants, and numbers of moral foundation labels, our method consistently outperforms the baseline methods. Comparing the results of base and base+, it is evident that incorporating MFs is beneficial, which aligns with findings from previous studies ziems2022moral; liu2025diagnosing.
Interestingly, for each model size, increasing the training dataset does not always yield performance gains. We believe one significant reason is that the additional data does not substantially expand the range of unique moral situations, and, as we discussed in the Section 1, moral reasoning depends on interpreting the implicatures of such situations taguchi2011teaching; taguchi2018contexts. On the other hand, for each dataset size, using larger models also does not always result in improved performance for both the baseline and our method. We believe this is because a training dataset of the same size is insufficient to modify the knowledge that larger LLMs have already acquired during pre-training compared to smaller models. A notable feature of our proposed pragmatic inference is that, unlike the baseline methods whose performance drops significantly as the number of underlying MFs increases, our method exhibits much smaller performance degradation.

### v3

@@ 1 Introduction
With the advancement of Large Language Models (LLMs), how to enhance their capability of moral reasoning in order to regulate their performance in accordance with human moral values has attracted scholarly interest. However, it still remains as a challenging task due to the gap between distributional semantics used by LLMs and the pragmatics-based nature of moral reasoning  Liu et al. (2025b); metaphorically speaking, the gap between what is said and what is morally implied. Distributional semantics follow Firth’s (1957) distributional hypothesis which argues that linguistic units have similar meaning if they occur in similar contexts. In contrast, moral reasoning goes beyond this structuralism view of language and requires the capability of deriving conclusions from implicature of language use, moral norms and behavioral consequences, and observations of discursive rules of conduct Spencer-Oatey and Xing (2019). These unsaid variables need to be linguistically manifested so that the gap between distributional semantics and pragmatics in moral reasoning can be closed, which increases the methodological generalizability that previous studies have thus far struggled to achieve Rao et al. (2023); Preniqi et al. (2024); Jiang et al. (2025).
@@
Table 1: Moral Reasoning Example From Ziems et al. (2022). The Moral Judgment for the given reply is Agree, and the Rule of Thumb (RoT) represents a judgment regarding an action mentioned in the prompt-reply, grounded in the moral foundation of fairness.
Table 1 demonstrates an example of the several elements that are involved in moral reasoning: rule of thumb (RoT), moral foundation, and moral judgment Ziems et al. (2022). Thus far, learning paradigms, which aim to equip LLMs with moral reasoning, generally train models to directly map a situation (Prompt and Reply) to moral judgment, with or without a reference to moral foundations and/or RoT  Hendrycks et al. (2020); Forbes et al. (2020); Jiang et al. (2021); Ziems et al. (2022); Jiang et al. (2025). Direct mapping often requires a huge amount of data to enable LLMs to capture the statistical correlations between situations and moral reasoning objectives.
On the other hand, we have witnessed the effectiveness of pragmatic inference used in other morality-relevant tasks, such as the identification of social biases and toxicity  Sap et al. (2020); Zhou et al. (2023); Chen and Wang (2025). For example,  Sap et al. (2020) developed a structured inference path that enables the identification of stereotypical implications encoded in explicit toxic language. Chen and Wang (2025) developed an inference chain that provides LLMs with the freedom to interpret online neologisms and rephrase the literal meanings of an input. Their method improved the performance of LLMs in recognizing highly implicit toxic language by 12% to 20%. These approaches are consonant with our goal of bridging the gap between distributional semantics and pragmatics, for instance, through interpreting literal meanings Chen and Wang (2025).
However, moral reasoning requires different inferential methods from the aforementioned toxicity or bias identification. Unlike toxic language (e.g., “lowering employment standards for females” implying “females are less capable”), the moral foundations***Please refer to Section 2 for more discussion on Moral Foundations Theory for moral reasoning. that underlie a moral situation are often not encoded by certain linguistic cues at all  Clark and Schaefer (1989); Graesser et al. (1994); Sap et al. (2022). Inferring them necessitates context-specific interpretations of abstract moral rules (e.g., the rule “do not hurt others” may manifest as “do not push another child to the floor” on a children’s playground and “do not emotionally abuse your partner” for adult relationships). Therefore, abstract moral foundations, rules of conduct, and context all need to be sufficiently textualized so as to be incorporated in machines’ reasoning process.
@@ 2 Related Works
Distributional Semantics is one of the theoretical foundations for embeddings. It interprets the meaning of a linguistic unit by additional signs (e.g., words, emoji) surrounding it Harris (1954); Firth (1957); McDonald and Ramscar (2001); Lenci and Sahlgren (2023); Boleda (2020). However, due to its nature as a structuralism approach to understanding language, distributional semantics suffers limitations. It treats the “context” purely as the texts provided, neglecting the many “unsaid” and latent parts of real interactions, such as the social constraints that a speaker encounters, the effects that a speech creates on the listeners, and the delivery of intention Lenci and others (2008); Purver and Sadrzadeh (2015). These are studied in pragmatics, where LLMs have shown several deficiencies, e.g., understanding implicature Cong (2024), comprehending metaphors and humours Ruis et al. (2023); Barattieri di SanPietro et al. (2023), and moral alignment Khamassi et al. (2024); Liu et al. (2024).
@@
Pragmatic Inference and Moral Reasoning. In this work, we leverage pragmatic inference to facilitate LLMs to achieve stronger performance in moral reasoning. Pragmatic inference refers to the process of deriving conclusions of meanings based on context, intention, and language use Elder (2024). The inferential process is often not verbally realized and, instead, involves evaluations made at the metapragmatic level Kádár and Haugh (2013). That is, humans construct meta links that connect an instance of language use to certain social qualities (e.g., linking the phrase “thank you” to the concept of politeness). They learn these metapragmatic links over their physical and social experiences, which are entirely absent in LLMs. These limitations suggest that: (i) the inferential process, that is not verbally realized and often does not need to be for humans, needs to be textualized so that LLMs can learn; and (ii) the context, where humans gain their experiences, needs to be sufficiently textualized, including both societal (e.g., social norms) and situational contexts (e.g., situated interpretations of language use). Without effective methods to textualize them, it is not a surprise that moral reasoning has long been a challenge in the broader field of artificial intelligence Allen et al. (2005); Van Rysewyk and Pontier (2014); Wallach et al. (2020).
@@
Moral Foundations Theory (MFT). Haidt and Joseph (2004) and Graham et al. (2013) describe a framework for understanding human morals and moral reasoning. MFT hypothesizes that moral judgments are guided by domain-specific foundations, i.e., care, fairness, loyalty, liberty, authority, and sanctity. These foundations capture the diversity of moral concerns and have been widely adopted in computational social science Johnson and Goldwasser (2018); Wendell and Tatalovich (2021); Reinig et al. (2024). They provide a principled way to annotate, analyze, and model moral dimensions in language, making them particularly relevant for studying how LLMs represent and reason about morality Forbes et al. (2020); Ziems et al. (2022).
@@ 3 Preliminary
Benchmark. In this paper, we adopt the Moral Integration Corpus (MIC) benchmark and its annotations from Ziems et al. (2022). We chose the MIC benchmark because it closely reflects how humans make moral judgments in daily communication and provides richer contextual information. One advantage of the MIC dataset is that it provides levels of agreement between the annotators, which allows us to focus on the less controversial part of their data. Table 1 presents an example from MIC. The example consists of a Prompt (“Do people actually shit themselves when they get very scared??”) and a Reply (“I don’t think I’ve ever heard of that happening, but I guess it’s possible..”), with three evaluations annotated: agreement with the reply as moral judgment, RoT, and MF. The MIC benchmark utilizes the six MFs defined by Graham et al. (2013): care, fairness, liberty, loyalty, authority, and sanctity. We take two of the evaluations as our objectives: moral judgment (agree, neutral, or disagree with the reply) and MFs. RoT, instead, is used as part of the learning paradigm. RoT is not selected as a learning objective because: (i) it does not have a fixed answer and, more crucially, multiple RoTs may exist in one situationWe refer readers to Section 3.1 of Ziems et al. (2022) for further details., and (ii) it has a rather dynamic relationship with MFs.
Experimental Settings. Following the steps of our pragmatic inference frameworks (detailed in Section 4), we design several questions and prompt the deepseek model to answer them. The step-wise answers are then used as the training dataThe data used in our experiments is available in the anonymized GitHub repository: https://anonymous.4open.science/r/moralpragmatics-F891. for fine-tuning the LLMs towards moral reasoning objectives. We chose Llama3-3.2-1B, Llama3-3.2-3B Grattafiori et al. (2024) and Mistral-7B (Mistral-7B-v0.3 Jiang et al. (2023)) as our backbone models, and leverage the supervised fine-tuning objective to train those models by following previous studies Ziems et al. (2022); Liu et al. (2025b). We intentionally select smaller models, as they are generally considered less capable than their larger counterparts. We aim to demonstrate the effectiveness of our pragmatic inference frameworks through the strong performance achieved by these smaller models.
@@
For baseline approaches, we follow prior work Forbes et al. (2020); Ziems et al. (2022) and consider two settings: (1) The base setting. For the moral judgment task, we compute $y_{J}=f_{\theta}(x_{S})$; and for MFC, we compute $y_{M}=f_{\theta}(x_{R})$. (2) The base+ setting. For the moral judgment task, we compute $y_{J}=f_{\theta}(x_{S},y_{M})$; and for MFC task, we compute $y_{M}=f_{\theta}(x_{R},d_{m})$. We additionally include the definitions of the six MFs ($d_{m}$) in the prompts. When the objective is moral judgment, we also explicitly provide the ground-truth MFs in the prompts. In other words, we intentionally provide the models with all available information in order to build strong baselines, which are then used to compare with the results of our framework. The setting of our proposed method is in line with the base+ setting but we incorporate an additional inference process in the input, e.g., $y_{J}=f_{\theta}(x_{S},y_{M},\{\text{inference}\})$.
@@ 4.1 Motivation
Proposing a pragmatic inference method for moral reasoning is challenging because there is no linguistic theory that clearly defines what constitutes an appropriate pragmatic inference for LLMs. Although several theories describe how humans make moral judgments, LLMs lack human-like linguistic competence Bender and Koller (2020); Mahowald et al. (2024), making these theories unsuitable for direct application. On the other hand, LLMs are well known to capture distributional knowledge (i.e., distributional semantics) through statistical associations, and prior works have shown that they cannot leverage this distributional knowledge for satisfactory moral reasoning performance Liu et al. (2025c); Lenci and Sahlgren (2023). However, the distributional knowledge is much beyond distributional semantics, this is because pre-training corpora also contains static and context-limited texts Graesser et al. (1994); Sap et al. (2022), forcing LLMs not only capture distributional semantics but also approximate pragmatic cues through shallow statistical correlations.
For instance, the statement “calling your uncle by his given name is rude” is acceptable in many East Asian cultures but may not be perceived the same way elsewhere. Because pre-training texts provide no explicit contextual cues about the cultural setting in which such statements are embedded, LLMs may learn this norm only by associating negative descriptors “is rude” with the action “calling your uncle by his name” even though they lack any actual cultural understanding Pawar et al. (2025). This implies that LLMs can leverage these superficial associations to predict certain pragmatic cues. However, the extent to which they can do so remains unknown, and we cannot rely on a purely semantics-driven approach to simplify the pragmatic inference. We hypothesize that MFT-driven pragmatic inference can help LLMs address the challenges outlined above, based on the following two perspectives.
@@
From the linguistic standpoint, our design of pragmatic inference methods is guided by three principles: (1) MFs are common ground which underlie all interactions. This has been supported by the theory and experiments of MFs Graham et al. (2013); Abdulhai et al. (2024). (2) MFs are abstract values extracted from discursive situations and hence need to be reified as context-specific interpretations Graham et al. (2011b). It is these context-specific interpretations of MFs which provide effective semantic cues for the models to reason morally. (3) Moral judgment is the judgment of good or bad by its nature, more specifically, the good/bad evaluation placed on the outcome of an action Nozick (1968); Schaich Borg et al. (2006); Emelin et al. (2021). Therefore, it is important to build the connections between contexts, discursive interpretations of MFs, and the outcome (i.e., behavioral consequence).
From a statistical standpoint, Liu et al. (2025b) demonstrate that the pragmatic gap between moral situations and moral reasoning objectives introduces greater degradation in language modeling performance than purely semantics-driven tasks. This is because LLMs tend to develop spurious associations, between moral situations and reasoning objectives, that undermine distributional semantics. To address this, MFT provides a principled signal: all moral situations can be interpreted through its lens. The key, then, is to leverage our pragmatic inference methods to supply linguistic cues that appropriately map moral situations to the LLMs’ hidden state space that is characterized by MFs, thereby alleviating the degradation of distributional semantics.
@@ 4.2 Moral Foundations Classification (MFC)
Figure 1 presents the prompting questions of the inferential steps used in the MFC task. The input here is a simple RoT and the definitions of six MFs. The inference steps are built as such: (1) Step 1 identifies the action encoded by the RoT Forbes et al. (2020). This step prepares the target (action) for moral evaluation. (2) Step 2 elicits the outcome of the action, namely, the behavioral consequence. One element of pragmatic inference is social norms, the variable that is probably most important in this task, given that RoTs were provided here without context. Norms for moral evaluations feature the physical and emotional harm to others that an action may incur Huebner et al. (2010); Kádár and Márquez-Reiter (2015); Turiel (1983). Therefore, Step 2 asks for the behavioral consequence as the linguistic manifestation of the social norms. This step setting also aligns with prior NLP studies showing that outcomes help establish the link between an action and its moral judgment Lindner et al. (2017); Lindner and Möllney (2019). (3) Step 3 connects the action and its consequence to relevant MFs by textualizing their metapragmatic links.
Table 2 reports performance on development set instances with 1–3 moral foundation labels. Moral situations with more than three labels are relatively rare. We also report the average performance for each experimental setting. Across all training dataset sizes, model variants, and numbers of moral foundation labels, our method consistently outperforms the baseline methods. Comparing the results of base and base+, it is evident that incorporating MFs is beneficial, which aligns with findings from previous studies Ziems et al. (2022); Liu et al. (2025b).
Interestingly, for each model size, increasing the training dataset does not always yield performance gains. We believe one significant reason is that the additional data does not substantially expand the range of unique moral situations, and, as we discussed in the Section 1, moral reasoning depends on interpreting the implicatures of such situations Taguchi (2011, 2018). On the other hand, for each dataset size, using larger models also does not always result in improved performance for both the baseline and our method. We believe this is because a training dataset of the same size is insufficient to modify the knowledge that larger LLMs have already acquired during pre-training compared to smaller models. A notable feature of our proposed pragmatic inference is that, unlike the baseline methods whose performance drops significantly as the number of underlying MFs increases, our method exhibits much smaller performance degradation.